OpenStudy (anonymous):

3x+4y+-10 and y=-3/4x-5/2

OpenStudy (anonymous):

somebody????????????????????? help

OpenStudy (mkmkasim):

what's your question?

OpenStudy (anonymous):

how i resolve this

OpenStudy (mkmkasim):

i mean do you need to solve for x and y ?

OpenStudy (anonymous):

3x+4y+-10 y=-3/4x-5/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please!!!!!!!

OpenStudy (mkmkasim):

ok let me see....

OpenStudy (anonymous):

nice

OpenStudy (mkmkasim):

well i guess x could equal 0, and y equal -5/2....

OpenStudy (mkmkasim):

if you plug them in they work.... i don't know why but it doesn't seem right...

OpenStudy (anonymous):

your x is correct but i dont know how to find y

OpenStudy (mkmkasim):

plug x=0 into y=-3/4x -5/2

OpenStudy (anonymous):

i pluged x=0 into the first one so it would be 3x+4y=-10 3(0) +4y=-10 -4y=-10-4 y=-14??

OpenStudy (mkmkasim):

ok so your algebra is wrong here. to make it easier btw...if you just plug 0 in for x in the second equation you get the answer of -5/2.....

OpenStudy (mkmkasim):

but if you want to use the first this is what you do

OpenStudy (mkmkasim):

3(0)+4y=-10 4y=-10 y=-10/4 or -5/2

OpenStudy (anonymous):

i need to find them in points that you can plot in a graph

OpenStudy (mkmkasim):

well if you know x=0, and y=-5/2.... your coordinate is (0,-5/2)

OpenStudy (anonymous):

okay thank you

OpenStudy (mkmkasim):

thanks!

OpenStudy (anonymous):

yup!